OpenStudy (anonymous):

I have a question about elimination in a system of equations. I'm in grade 10 and understand elimination, but my answer includes a fraction, and I still have trouble with fractions in an equation. The equations are 4x-6y=10 (minus) 4x+y=8 which is -7y=2. I'm having trouble isolating the y.

OpenStudy (mertsj):

divide both sides of your equation by -7

OpenStudy (anonymous):

I did that, but its a decimal, and that gets messy

OpenStudy (mertsj):

leave it in fractional form. y=-2/7

OpenStudy (mertsj):

\[4x+y=8\] \[4x-\frac{2}{7}=8\] \[4x=8 \frac{2}{7}=\frac{58}{7}\]

OpenStudy (mertsj):

Now multiply both sides by 1/4

OpenStudy (anonymous):

Full original question... Consider the linear system 2x - 3y = 5 and 4x + y = 8. Solve by elimination. So, I did this. 2x-3y=5(x2) so... 4x-6y=10 4x+y=8................4x+y=8 So then, I subtract those two equation (on the right) And I get -7y=2 Now what?

OpenStudy (mertsj):

You already found y. Why are you doing it over?

OpenStudy (anonymous):

no, i have to isolate the y

OpenStudy (mertsj):



OpenStudy (mertsj):